A rectangle is 6 times as long as it is wide. The area is 96 square feet. What is the perimeter of the rectangle?

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

A rectangle is 6 times as long as it is wide. The area is 96 square feet. What is the perimeter of the rectangle?
-----
L = 6W
L*W = 96
---
6W*W = 96
W*W = 16
W = 4 ft
------------
Perimeter = 2W + 2L

Answer by stanbon(75887)

A rectangle is 6 times as long as it is wide. The area is 96 square feet. What is the perimeter of the rectangle?
---------
width:: x
length:: 6x
----
Equation:
x*6x = 96
----
6x^2 = 96
----
x^2 = 16
---
x = 4 (width)
6x = 24 (length)
---------------
Perimeter = 2(length+width)
P = 2(24+4)
P = 2(28)
P = 56
-----------
Cheers,
Stan H.
-----------